# Quantum conditional entropies and fully entangled fraction of states with maximally mixed marginals

### v1

@@ VIII Conclusion
The present work discusses the relationship between conditional entropies and fully entangled fraction, as they provide important yardsticks for the merit of quantum states in information processing tasks. In $2\otimes 2$ system, for Werner state, we obtain the upper bounds of conditional von Neumann entropy provided its fully entangled fraction is greater than half. Also, the lower bound of the fully entangled fraction is obtained in terms of eigenvalue if its conditional Rényi $\alpha-$entropy is negative. For a two-qubit Weyl state, the negativity of conditional Rényi $2-$entropy is a sufficient condition for being useful for teleportation, and the upper bounds to the conditional Rényi $2-$entropy is obtained whenever it is useful for teleportation. Further, extending our studies to arbitrary $d\otimes d$ systems, we observe that the fully entangled fraction of an isotropic state is lower bounded provided its conditional von Neumann entropy is negative. We study the relation between fully entangled fraction and conditional entropies for for two qudit Werner states, rank deficient states and generalized Bell diagonal states. Also, we investigate the relation of $k-$copy steerability, $k-$copy nonlocality and conditional entropies. If the conditional von Neumann entropy of two-qubit isotropic is negative then it is $k-$copy steerable, same behaviour holds in the case of conditional Rényi $2-$entropy. On considering a few copies ($k=2$) of isotropic state in six dimensions, the negativity of conditional von Neumann entropy and conditional Rényi $2-$entropy guarantees the $k-$copy steerability. The negativity of conditional Rényi $2-$entropy of two-qubit Weyl state gives the sufficient condition for $k-$copy nonlocality. The same holds for some non-Weyl states. We show that the negativity of conditional Rényi $2-$entropy of state (mixture of arbitrary pure state and maximally mixed noise) for dimension $d=3,4,5$ is the sufficient condition for $k-$copy nonlocality, and the same holds whenever the conditional von Neumann entropy is negative. Further, we discuss the application of fully entangled fraction in the context of work extraction.
@@
We observe that in general the relation between FEF (particularly when the threshold value $FEF>1/d$) and negativity of conditional entropy is intricate. This calls for further probe for general quantum states. As we observe, such relations are significantly linked with quantum information processing protocols and thus further investigation will shed light on the efficacy of quantum states in information processing tasks. The implications of this study also carry significance in foundational studies like $k-$ copy steerability and $k-$ copy nonlocality and thus calls for further attention.
Acknowledgement
@@
NG acknowledges support from SERB(DST) MATRICS grant vide file number MTR/2022/000101.

### v2

@@ VIII Conclusion
The present work discusses the relationship between conditional entropies and fully entangled fraction, as they provide important yardsticks for the merit of quantum states in information processing tasks. In $2\otimes 2$ system, for Werner state, we obtain the upper bounds of conditional von Neumann entropy provided its fully entangled fraction is greater than half. Also, the lower bound of the fully entangled fraction is obtained in terms of eigenvalue if its conditional Rényi $\alpha-$entropy is negative. For a two-qubit Weyl state, the negativity of conditional Rényi $2-$entropy is a sufficient condition for being useful for teleportation, and the upper bounds to the conditional Rényi $2-$entropy is obtained whenever it is useful for teleportation. Further, extending our studies to arbitrary $d\otimes d$ systems, we observe that the fully entangled fraction of an isotropic state is lower bounded provided its conditional von Neumann entropy is negative. We study the relation between fully entangled fraction and conditional entropies for for two qudit Werner states, rank deficient states and generalized Bell diagonal states. Also, we investigate the relation of $k-$copy steerability, $k-$copy nonlocality and conditional entropies. If the conditional von Neumann entropy of two-qubit isotropic is negative then it is $k-$copy steerable, same behaviour holds in the case of conditional Rényi $2-$entropy. On considering a few copies ($k=2$) of isotropic state in six dimensions, the negativity of conditional von Neumann entropy and conditional Rényi $2-$entropy guarantees the $k-$copy steerability. The negativity of conditional Rényi $2-$entropy of two-qubit Weyl state gives the sufficient condition for $k-$copy nonlocality. The same holds for some non-Weyl states. We show that the negativity of conditional Rényi $2-$entropy of state (mixture of arbitrary pure state and maximally mixed noise) for dimension $d=3,4,5$ is the sufficient condition for $k-$copy nonlocality, and the same holds whenever the conditional von Neumann entropy is negative. Further, we discuss the implications of the relations obtained in the context of work extraction, faithful entanglement, and entropic uncertainty relation.
@@
We observe that in general the relation between FEF (particularly when the threshold value $\text{FEF}>1/d$) and negativity of conditional entropy is intricate. This calls for further probe for general quantum states. As we observe, such relations are significantly linked with quantum information processing protocols and thus further investigation will shed light on the efficacy of quantum states in information processing tasks. The implications of this study also carry significance in foundational studies like $k-$ copy steerability and $k-$ copy nonlocality and thus calls for further attention.
@@
Acknowledgement: NG acknowledges support from SERB-DST (India) MATRICS grant vide file number MTR/2022/000101.
Declaration
All the authors contributed equally to the manuscript.
Data availability: Data sharing was not applicable to this article as no data sets were generated or analyzed during the current study.
Competing interests: The authors have no competing interests to declare. All co-authors have seen and agree with the contents of the manuscript, and there is no financial interest to report.